# Wireless Communication for Low-Altitude Economy with UAV Swarm Enabled Two-Level Movable Antenna System

### v1

@@ IV-A Single UE and Two UEs
For the scenario with two UEs, the resulting SINR/SNR of UE $k$ with the MRC, ZF, and MMSE beamforming can be similarly obtained by substituting $L$ with $LM$ in (14)-(16), respectively. Besides, the channel’s squared-correlation coefficient between UEs $k$ and $k^{\prime}$ is
${\xi_{\_}{k,k^{\prime}}}$ $=\frac{1}{{{L^{2}}{M^{2}}}}{\left|{\sum\limits_{\_}{l=1}^{L}{{\bf{a}}_{\_}{k,l}^{H}{{\bf{a}}_{\_}{k^{\prime},l}}}}\right|^{2}}$ (41)
$=\frac{1}{{{L^{2}}{M^{2}}}}{\left|{\sum\limits_{\_}{l=1}^{L}{{e^{-{\rm{j}}\frac{{2\pi}}{\lambda}\left({{\bm{\kappa}}_{\_}k^{T}-{\bm{\kappa}}_{\_}{k^{\prime}}^{T}}\right){{\bf{q}}_{\_}l}}}\sum\limits_{\_}{m=1}^{M}{{e^{-{\rm{j}}\frac{{2\pi}}{\lambda}{b_{\_}{l,m}}\Delta{\Phi_{\_}{k,k^{\prime}}}}}}}}\right|^{2}},$
where $\Delta{\Phi_{\_}{k,k^{\prime}}}\triangleq{\Phi_{\_}k}-{\Phi_{\_}{k^{\prime}}}$. It is observed from (41) that when _$\sum\nolimits_{\_}{m=1}^{M}{{e^{-{\rm{j}}\frac{{2\pi}}{\lambda}{b_{\_}{l,m}}\Delta{\Phi_{\_}{k,k^{\prime}}}}}}=0$, $\forall l$, we have ${\xi_{\_}{k,k^{\prime}}}=0$.
Similar to Theorem 1, an optimal solution of $\left\{{{b_{\_}{l,m}}}\right\}$ to (41) that achieves zero objective value is
${b_{\_}{l,m}}=\left({m-1}\right)\frac{{\left({\zeta_{\_}l^{\star}+1/M}\right)\lambda}}{\left|{\Delta{\Phi_{\_}{k,k^{\prime}}}}\right|},$ (42)
where the common coefficient ${\zeta_{\_}l^{\star}}$ is derived to satisfy the mutual coupling avoidance constraint between adjacent MA elements for each UAV, given by
${\zeta_{\_}l^{\star}}=\left\{\begin{split}&\frac{{{{\tilde{\nu}}_{\_}{\min}}}}{M},\ \ \ \ \ \ {\rm if}\ \bmod\left({{{\tilde{\nu}}_{\_}{\min}}+1,M}\right)\neq 0,\\ &\frac{{{{\tilde{\nu}}_{\_}{\min}}+1}}{M},\ {\rm otherwise}.\end{split}\right.$ (43)
with ${{\tilde{\nu}}_{\_}{\min}}=\left\lceil{M{{\tilde{d}}_{\_}{\min}}\left|{\Delta{\Phi_{\_}{k,k^{\prime}}}}\right|/\lambda-1}\right\rceil$ common to all UAVs. Moreover, MA position ${b_{\_}{l,m}}\in\tilde{\cal C}$, $\forall l,m$, can be satisfied when $\left|{{b_{\_}{l,M}}-{b_{\_}{l,1}}}\right|=\left({M-1}\right)\frac{{\left({\zeta_{\_}l^{\star}+1/M}\right)\lambda}}{{\left|{\Delta{\Phi_{\_}{k,k^{\prime}}}}\right|}}$ is no greater than the maximum distance between any two points in ${\tilde{\cal C}}$.
@@
$\mathop{\min}\limits_{\_}{k\in{\cal K}}{{{\log}_{\_}2}\left({1+{{\bar{P}}_{\_}k}{{\left|{{\alpha_{\_}k}}\right|}^{2}}LM}\right)}.$ (44)
@@
In fact, the results in (42)-(43) correspond to the case where all MA arrays have the identical architecture. To gain further insights, we express ${{\bf{b}}_{\_}{l,m}}$ as ${{\bf{b}}_{\_}{l,m}}={{\bf{b}}_{\_}m}={\left[{{b_{\_}m},0,0}\right]^{T}}$, $\forall l$, the channels’ squared-correlation coefficient in (41) is reduced to
@@
${\xi_{\_}{k,k^{\prime}}}=\frac{1}{{{L^{2}}}}{\left|{\sum\limits_{\_}{l=1}^{L}{{e^{-{\rm{j}}\frac{{2\pi}}{\lambda}\left({{\bm{\kappa}}_{\_}k^{T}-{\bm{\kappa}}_{\_}{k^{\prime}}^{T}}\right){{\bf{q}}_{\_}l}}}}}\right|^{2}}\underbrace{\frac{1}{{{M^{2}}}}{{\left|{\sum\limits_{\_}{m=1}^{M}{{e^{-{\rm{j}}\frac{{2\pi}}{\lambda}{b_{\_}{m}}\Delta{\Phi_{\_}{k,k^{\prime}}}}}}}\right|}^{2}}}_{\_}{\leq 1}.$ (45)
@@
It is observed that (45) differs from (19) in the term ${\frac{1}{{{M^{2}}}}{{|{\sum\nolimits_{\_}{m=1}^{M}{{e^{-{\rm{j}}\frac{{2\pi}}{\lambda}{b_{\_}{m}}\Delta{\Phi_{\_}{k,k^{\prime}}}}}}}|}^{2}}}\leq 1$, which characterizes the capability of the MA array mounted on each UAV to distinguish UE $k$ and $k^{\prime}$ in the spatial domain. Thus, the IUI can be mitigated by adjusting not only the UAV swarm placement positions, but also the local positions of MAs at all UAVs.
@@
In particular, when the USA is adopted, i.e., ${b_{\_}m}=\left({m-1}\right)\Gamma d=\left({m-1}\right)\Gamma\lambda/2$, where $\Gamma\geq 1$ denotes the sparsity level [44, 26], (45) reduces to
@@
${\xi_{\_}{k,k^{\prime}}}$ $=\frac{1}{{{L^{2}}}}{\left|{\sum\limits_{\_}{l=1}^{L}{{e^{-{\rm{j}}\frac{{2\pi}}{\lambda}\left({{\bm{\kappa}}_{\_}k^{T}-{\bm{\kappa}}_{\_}{k^{\prime}}^{T}}\right){{\bf{q}}_{\_}l}}}}}\right|^{2}}\times$ (46)
$\ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \frac{1}{{{M^{2}}}}\left|{\frac{{\sin\left({\frac{{\pi M\Gamma\Delta{\Phi_{\_}{k,k^{\prime}}}}}{2}}\right)}}{{\sin\left({\frac{{\pi\Gamma\Delta{\Phi_{\_}{k,k^{\prime}}}}}{2}}\right)}}}\right|^{2}.$
Such a result resembles the existing modular array [45], while a subtle difference lies in that modular array is composed of multiple compact arrays where antenna spacing is separated by half wavelength, rather than USAs. Then, the coefficient ${\xi_{\_}{k,k^{\prime}}}$ is equal to zero when sparsity level satisfies
@@
$\Gamma\in\left\{{\left.{\frac{{2\varepsilon}}{{M\Delta{\Phi_{\_}{k,k^{\prime}}}}}}\right|\varepsilon\in{\mathbb{Z}},{\frac{{2\varepsilon}}{{M\Delta{\Phi_{\_}{k,k^{\prime}}}}}\geq 1},\bmod\left({\varepsilon,M}\right)\neq 0}\right\}.$ (47)
@@
Figure 2: The channels’ squared-correlation coefficient ${\xi_{\_}{1,2}}$ versus $\varsigma$ and $\Gamma$.
@@
Fig. 2 shows the channels’ squared-correlation coefficient ${\xi_{\_}{1,2}}$ versus $\varsigma$ and $\Gamma$ for the two-UE communication, by considering the placement positions given in (20) and (21). The UE directions are $\left({{\theta_{\_}1},{\phi_{\_}1}}\right)=\left({{{60}^{\circ}},{{30}^{\circ}}}\right)$ and $\left({{\theta_{\_}2},{\phi_{\_}2}}\right)=\left({{{30}^{\circ}},{{0}^{\circ}}}\right)$, respectively. The number of UAVs is $L=4$, and each UAV is equipped with a USA with $M=4$ antennas. For convenience of presentation, the coefficient ${\xi_{\_}{1,2}}$ below $-200$ dB is truncated to $-200$ dB. It is observed that when $\varsigma$ satisfies (54) in Appendix A, the coefficient ${\xi_{\_}{1,2}}$ is equal to zero, i.e., the channel of UE 1 is orthogonal to that of UE 2, and an IUI-free communication can be obtained. It is also observed that the coefficient ${\xi_{\_}{1,2}}$ can be reduced by adjusting the sparsity level $\Gamma$, and thus providing an extra DoF for IUI mitigation as compared to the single-antenna UAV.
@@ IV-B1 Optimization of ${\bf V}$ With Given ${\bf Q}$ and ${\bf B}$
IV-B2 Optimization of ${{\bf{q}}_{\_}l}$ With Given ${\bf V}$, $\left\{{{\bf{q}}_{\_}{l^{\prime}},\forall l^{\prime}\neq l}\right\}$, and $\bf B$
@@
Let ${{{\bf{\tilde{h}}}}_{\_}{i,l}}\in{{\mathbb{C}}^{\left({L-1}\right)M\times 1}}$ denote the resulting channel after removing ${{\bf{h}}_{\_}{i,l}}\in{{\mathbb{C}}^{M\times 1}}={\alpha_{\_}i}{{\bf{a}}_{\_}{i,l}}$ from ${{\bf{h}}_{\_}i}$, ${{{\bf{\tilde{v}}}}_{\_}{k,l}}\in{{\mathbb{C}}^{\left({L-1}\right)M\times 1}}$ denote the resulting beamforming after removing ${{\bf{v}}_{\_}{k,l}}\in{\mathbb{C}}^{M\times 1}$ from ${{\bf{v}}_{\_}k}$, where ${{\bf{h}}_{\_}{i,l}}$ and ${{\bf{v}}_{\_}{k,l}}$ denote the $l$-th block of ${{\bf{h}}_{\_}i}$ and ${{\bf{v}}_{\_}k}$, respectively. For the case of multi-antenna UAV, after some manipulations, ${f_{\_}{k,i}}={\left|{{\bf{v}}_{\_}k^{H}{{\bf{h}}_{\_}i}}\right|^{2}}$ can be expressed as
@@
${f_{\_}{k,i}}$ =_$=\underbrace{\sum\limits_{\_}{l^{\prime}=1,l^{\prime}\neq l}^{L}{2{{\left|{{\alpha_{\_}i}}\right|}^{2}}{\rm{Re}}\left\{{{e^{-{\rm{j}}\frac{{2\pi}}{\lambda}{\bm{\kappa}}_{\_}i^{T}\left({{{\bf{q}}_{\_}l}-{{\bf{q}}_{\_}{l^{\prime}}}}\right)}}{\bf{d}}_{\_}{i,l}^{H}{{\bf{V}}_{\_}{k,l,l^{\prime}}}{{\bf{d}}_{\_}{i,l^{\prime}}}}\right\}}}_{\_}{{{\tilde{g}}_{\_}{k,i,l}}}$ (48)
$+\underbrace{{{\left|{{\alpha_{\_}i}}\right|}^{2}}{\bf{d}}_{\_}{i,l}^{H}{{\bf{V}}_{\_}{k,l,l}}{{\bf{d}}_{\_}{i,l}}+{\bf{\tilde{h}}}_{\_}{i,l}^{H}{{{\bf{\tilde{V}}}}_{\_}{k,l,l}}{{{\bf{\tilde{h}}}}_{\_}{i,l}}}_{\_}{{{\hat{g}}_{\_}{k,i,l}}},$
@@
where ${{\bf{d}}_{\_}{i,l}}\in{{\mathbb{C}}^{M\times 1}}\triangleq{[{e^{{\rm{j}}\frac{{2\pi}}{\lambda}{\bm{\kappa}}_{\_}i^{T}\left({{{\bf{b}}_{\_}{l,1}}-{{\bf{q}}_{\_}r}}\right)}},\cdots,}$ ${{e^{{\rm{j}}\frac{{2\pi}}{\lambda}{\bm{\kappa}}_{\_}i^{T}\left({{{\bf{b}}_{\_}{l,M}}-{{\bf{q}}_{\_}r}}\right)}}]^{T}}$, ${{\bf{V}}_{\_}{k,l,l^{\prime}}}\in{{\mathbb{C}}^{M\times M}}\triangleq{{\bf{v}}_{\_}{k,l}}{\bf{v}}_{\_}{k,l^{\prime}}^{H}$, and ${{{\bf{\tilde{V}}}}_{\_}{k,l,l}}\in{{\mathbb{C}}^{\left({L-1}\right)M\times\left({L-1}\right)M}}\triangleq{{{\bf{\tilde{v}}}}_{\_}{k,l}}{\bf{\tilde{v}}}_{\_}{k,l}^{H}$. Moreover, ${{{\hat{g}}_{\_}{k,i,l}}}$ is a constant term, and ${{{\tilde{g}}_{\_}{k,i,l}}}$ can be expressed in terms of ${{{\bf{q}}_{\_}l}}$ as
${{\tilde{g}}_{\_}{k,i,l}}$ =_$=\sum\limits_{\_}{l^{\prime}=1,l^{\prime}\neq l}^{L}{2{{\left|{{\alpha_{\_}i}}\right|}^{2}}\left|{{\bf{d}}_{\_}{i,l}^{H}{{\bf{V}}_{\_}{k,l,l^{\prime}}}{{\bf{d}}_{\_}{i,l^{\prime}}}}\right|\times}$ (49)
$\ \ {\rm{cos}}\left({\frac{{2\pi}}{\lambda}{\bm{\kappa}}_{\_}i^{T}\left({{{\bf{q}}_{\_}l}-{{\bf{q}}_{\_}{l^{\prime}}}}\right)-\angle{\bf{d}}_{\_}{i,l}^{H}{{\bf{V}}_{\_}{k,l,l^{\prime}}}{{\bf{d}}_{\_}{i,l^{\prime}}}}\right).$
@@
A closer look at (49) shows that it has a similar form to (32), and its lower and upper bounds can be correspondingly obtained. As a result, the optimization of ${\bf{q}}_{\_}l$ for the multi-antenna UAV can follow the similar procedure as in (39).
IV-B3 Optimization of ${b_{\_}{l,m}}$ With Given ${\bf V}$, ${\bf Q}$, and $\left\{{{b_{\_}{l^{\prime},m^{\prime}}},\forall l^{\prime}\neq l,m^{\prime}\neq m}\right\}$
The sub-problem of (P1) for optimizing ${b_{\_}{l,m}}$ is
$\mathop{\max}\limits_{\_}{{b_{\_}{l,m}}}$ $\ \mathop{\min}\limits_{\_}{k\in{\cal K}}\ {{R_{\_}k}}$ (50)
${\rm{s.t.}}$ $\ {b_{\_}{l,m}}\in{\tilde{\cal C}},$
$\ {\left|{{b_{\_}{l,m}}-{b_{\_}{l,m^{\prime}}}}\right|^{2}}\geq\tilde{d}_{\_}{\min}^{2},\ \forall m^{\prime}\neq m.$

### v2

@@ IV-A Single UE and Two UEs
For the scenario with two UEs, the resulting SINR/SNR of UE $k$ with the MRC, ZF, and MMSE beamforming can be similarly obtained by substituting $L$ with $LM$ in (15)-(17), respectively. Besides, when UPW model holds, the channel’s squared-correlation coefficient between UEs $k$ and $k^{\prime}$ is
${\xi_{k,k^{\prime}}}$ $=\frac{1}{{{L^{2}}{M^{2}}}}{\left|{\sum\limits_{l=1}^{L}{{\bf{a}}_{k,l}^{H}{{\bf{a}}_{k^{\prime},l}}}}\right|^{2}}$ (40)
$=\frac{1}{{{L^{2}}{M^{2}}}}{\left|{\sum\limits_{l=1}^{L}{{e^{{\rm{j}}\frac{{2\pi}}{\lambda}\left({{\bm{\kappa}}_{k}^{T}-{\bm{\kappa}}_{k^{\prime}}^{T}}\right){{\bf{q}}_{l}}}}\sum\limits_{m=1}^{M}{{e^{{\rm{j}}\frac{{2\pi}}{\lambda}{b_{l,m}}\Delta{\Phi_{k,k^{\prime}}}}}}}}\right|^{2}},$
where $\Delta{\Phi_{k,k^{\prime}}}\triangleq{\Phi_{k}}-{\Phi_{k^{\prime}}}$. It is observed from (40) that when $\sum\nolimits_{m=1}^{M}{{e^{{\rm{j}}\frac{{2\pi}}{\lambda}{b_{l,m}}\Delta{\Phi_{k,k^{\prime}}}}}}=0$, $\forall l$, we have ${\xi_{k,k^{\prime}}}=0$.
Similar to Theorem 1, an optimal solution of $\left\{{{b_{l,m}}}\right\}$ to (40) that achieves zero objective value is
${b_{l,m}}=\left({m-1}\right)\frac{{\left({\zeta_{l}^{\star}+1/M}\right)\lambda}}{\left|{\Delta{\Phi_{k,k^{\prime}}}}\right|},$ (41)
where the common coefficient $\zeta_{l}^{\star}$} is derived to satisfy the mutual coupling avoidance constraint between adjacent MA elements for each UAV, given by
${\zeta_{l}^{\star}}=\left\{\begin{split}&\frac{{{{\tilde{\nu}}_{\min}}}}{M},\ \ \ \ \ \ {\rm if}\ \bmod\left({{{\tilde{\nu}}_{\min}}+1,M}\right)\neq 0,\\ &\frac{{{{\tilde{\nu}}_{\min}}+1}}{M},\ {\rm otherwise}.\end{split}\right$. (42)
with ${{\tilde{\nu}}_{\min}}=\left\lceil{M{{\tilde{d}}_{\min}}\left|{\Delta{\Phi_{k,k^{\prime}}}}\right|/\lambda-1}\right\rceil$ common to all UAVs. Moreover, MA position ${b_{l,m}}\in\tilde{\cal C}$, $\forall l,m$, can be satisfied when $\left|{{b_{l,M}}-{b_{l,1}}}\right|=\left({M-1}\right)\frac{{\left({\zeta_{l}^{\star}+1/M}\right)\lambda}}{{\left|{\Delta{\Phi_{k,k^{\prime}}}}\right|}}$ is no greater than the maximum distance between any two points in $\tilde{\cal C}$}.
@@
$\mathop{\min}\limits_{k\in{\cal K}}{{{\log}_{2}}\left({1+{{\bar{P}}_{k}}{{\left|{{\alpha_{k}}}\right|}^{2}}LM}\right)}.$ (43)
@@
In fact, the results in (41)-(42) correspond to the case where all MA arrays have the identical architecture. To gain further insights, we express ${\bf{b}}_{l,m}$} as ${{\bf{b}}_{l,m}}={{\bf{b}}_{m}}={\left[{{b_{m}},0,0}\right]^{T}}$, $\forall l$, the channels’ squared-correlation coefficient in (40) is reduced to
@@
${\xi_{k,k^{\prime}}}=\frac{1}{{{L^{2}}}}{\left|{\sum\limits_{l=1}^{L}{{e^{{\rm{j}}\frac{{2\pi}}{\lambda}\left({{\bm{\kappa}}_{k}^{T}-{\bm{\kappa}}_{k^{\prime}}^{T}}\right){{\bf{q}}_{l}}}}}}\right|^{2}}\underbrace{\frac{1}{{{M^{2}}}}{{\left|{\sum\limits_{m=1}^{M}{{e^{{\rm{j}}\frac{{2\pi}}{\lambda}{b_{m}}\Delta{\Phi_{k,k^{\prime}}}}}}}\right|}^{2}}}_{\leq 1}.$ (44)
@@
It is observed that (44) differs from (20) in the term ${\frac{1}{{{M^{2}}}}{{|{\sum\nolimits_{m=1}^{M}{{e^{{\rm{j}}\frac{{2\pi}}{\lambda}{b_{m}}\Delta{\Phi_{k,k^{\prime}}}}}}}|}^{2}}}\leq 1$, which characterizes the capability of the MA array mounted on each UAV to distinguish UE $k$ and $k^{\prime}$ in the spatial domain. Thus, the IUI can be mitigated by adjusting not only the UAV swarm placement positions, but also the local positions of MAs at all UAVs.
@@
In particular, when the USA is adopted, i.e., ${b_{m}}=\left({m-1}\right)\Gamma d=\left({m-1}\right)\Gamma\lambda/2$, where $\Gamma\geq 1$ denotes the sparsity level [44, 26], (44) reduces to
@@
${\xi_{k,k^{\prime}}}$ $=\frac{1}{{{L^{2}}}}{\left|{\sum\limits_{l=1}^{L}{{e^{{\rm{j}}\frac{{2\pi}}{\lambda}\left({{\bm{\kappa}}_{k}^{T}-{\bm{\kappa}}_{k^{\prime}}^{T}}\right){{\bf{q}}_{l}}}}}}\right|^{2}}\times$ (45)
$\ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \frac{1}{{{M^{2}}}}\left|{\frac{{\sin\left({\frac{{\pi M\Gamma\Delta{\Phi_{k,k^{\prime}}}}}{2}}\right)}}{{\sin\left({\frac{{\pi\Gamma\Delta{\Phi_{k,k^{\prime}}}}}{2}}\right)}}}\right|^{2}.$
Then, the coefficient $\xi_{k,k^{\prime}}$} is equal to zero when sparsity level satisfies
@@
$\Gamma\in\left\{{\left.{\frac{{2\varepsilon}}{{M\Delta{\Phi_{k,k^{\prime}}}}}}\right|\varepsilon\in{\mathbb{Z}},{\frac{{2\varepsilon}}{{M\Delta{\Phi_{k,k^{\prime}}}}}\geq 1},\bmod\left({\varepsilon,M}\right)\neq 0}\right\}.$ (46)
@@
Figure 2: The channels’ squared-correlation coefficient $\xi_{1,2}$} versus $\varsigma$ and $\Gamma$.
@@
Fig. 2 shows the channels’ squared-correlation coefficient $\xi_{1,2}$} versus $\varsigma$ and $\Gamma$ for the two-UE communication, by considering the placement positions given in (21) and (22). The UE directions are $\left({{\theta_{1}},{\phi_{1}}}\right)=\left({{{30}^{\circ}},{{60}^{\circ}}}\right)$ and $\left({{\theta_{2}},{\phi_{2}}}\right)=\left({{{0}^{\circ}},{{0}^{\circ}}}\right)$, respectively. The number of UAVs is $L=4$, and each UAV is equipped with a USA with $M=4$ antennas. For convenience of presentation, the coefficient $\xi_{1,2}$} below 200-200 dB is truncated to 200-200 dB. It is observed that when $\varsigma$ satisfies (53) in Appendix A, the coefficient $\xi_{1,2}$} is equal to zero, i.e., the channel of UE 1 is orthogonal to that of UE 2, and an IUI-free communication can be obtained. It is also observed that the coefficient $\xi_{1,2}$} can be reduced by adjusting the sparsity level $\Gamma$, and thus providing an extra DoF for IUI mitigation as compared to the single-antenna UAV.
@@ IV-B1 Optimization of $\bf V$} With Given $\bf Q$} and $\bf B$}
IV-B2 Optimization of ${\bf{q}}_{l}$} With Given $\bf V$}, $\left\{{{\bf{q}}_{l^{\prime}},\forall l^{\prime}\neq l}\right\}$, and $\bf B$
@@
Let ${{{\bf{\tilde{h}}}}_{i,l}}\in{{\mathbb{C}}^{\left({L-1}\right)M\times 1}}$ denote the resulting channel after removing ${{\bf{h}}_{i,l}}\in{{\mathbb{C}}^{M\times 1}}={\alpha_{i}}{{\bf{a}}_{i,l}}$ from ${\bf{h}}_{i}$}, ${{{\bf{\tilde{v}}}}_{k,l}}\in{{\mathbb{C}}^{\left({L-1}\right)M\times 1}}$ denote the resulting beamforming after removing ${{\bf{v}}_{k,l}}\in{\mathbb{C}}^{M\times 1}$ from ${\bf{v}}_{k}$}, where ${\bf{h}}_{i,l}$} and ${\bf{v}}_{k,l}$} denote the $l$-th block of ${\bf{h}}_{i}$} and ${\bf{v}}_{k}$}, respectively. For the case of multi-antenna UAV, after some manipulations, ${f_{k,i}}={\left|{{\bf{v}}_{k}^{H}{{\bf{h}}_{i}}}\right|^{2}}$ can be expressed as
@@
${f_{k,i}}$ $=\underbrace{\sum\limits_{l^{\prime}=1,l^{\prime}\neq l}^{L}{2{{\left|{{\alpha_{i}}}\right|}^{2}}{\mathop{\rm Re}\nolimits}\left\{{{e^{{\rm{j}}\frac{{2\pi}}{\lambda}\left({{r_{i,l}}-{r_{i,l^{\prime}}}}\right)}}{\bf{d}}_{i,l}^{H}{{\bf{V}}_{k,l,l^{\prime}}}{{\bf{d}}_{i,l^{\prime}}}}\right\}}}_{{{\tilde{g}}_{k,i,l}}}$ (47)
$+\underbrace{{{\left|{{\alpha_{i}}}\right|}^{2}}{\bf{d}}_{i,l}^{H}{{\bf{V}}_{k,l,l}}{{\bf{d}}_{i,l}}+{\bf{\tilde{h}}}_{i,l}^{H}{{{\bf{\tilde{V}}}}_{k,l,l}}{{{\bf{\tilde{h}}}}_{i,l}}}_{{{\hat{g}}_{k,i,l}}},$
@@
where ${{\bf{d}}_{i,l}}={[{e^{-{\rm{j}}\frac{{2\pi}}{\lambda}{\bm{\kappa}}_{k,l}^{T}{{\bf{b}}_{l,1}}}},\cdots,{e^{-{\rm{j}}\frac{{2\pi}}{\lambda}{\bm{\kappa}}_{k,l}^{T}{{\bf{b}}_{l,M}}}}]^{T}}$, ${{\bf{V}}_{k,l,l^{\prime}}}\in{{\mathbb{C}}^{M\times M}}\triangleq{{\bf{v}}_{k,l}}{\bf{v}}_{k,l^{\prime}}^{H}$, and ${{{\bf{\tilde{V}}}}_{k,l,l}}\in{{\mathbb{C}}^{\left({L-1}\right)M\times\left({L-1}\right)M}}\triangleq{{{\bf{\tilde{v}}}}_{k,l}}{\bf{\tilde{v}}}_{k,l}^{H}$. Moreover, ${\hat{g}}_{k,i,l}$}} is a constant term, and ${\tilde{g}}_{k,i,l}$}} can be expressed in terms of ${\bf{q}}_{l}$}} as
${{\tilde{g}}_{k,i,l}}$ $=\sum\limits_{l^{\prime}=1,l^{\prime}\neq l}^{L}{2{{\left|{{\alpha_{i}}}\right|}^{2}}\left|{{\bf{d}}_{i,l}^{H}{{\bf{V}}_{k,l,l^{\prime}}}{{\bf{d}}_{i,l^{\prime}}}}\right|\times}$ (48)
$\ \ {\cos\left({\frac{{2\pi}}{\lambda}\left({{r_{i,l}}-{r_{i,l^{\prime}}}}\right)+\angle{\bf{d}}_{i,l}^{H}{{\bf{V}}_{k,l,l^{\prime}}}{{\bf{d}}_{i,l^{\prime}}}}\right)}.$
@@
A closer look at (48) shows that it has a similar form to (31), and its lower and upper bounds can be correspondingly obtained. As a result, the optimization of ${\bf{q}}_{l}$ for the multi-antenna UAV can follow the similar procedure as in (38).
IV-B3 Optimization of $b_{l,m}$} With Given $\bf V$}, $\bf Q$}, and $\left\{{{b_{l^{\prime},m^{\prime}}},\forall l^{\prime}\neq l,m^{\prime}\neq m}\right\}$
The sub-problem of (P1) for optimizing $b_{l,m}$} is
$\mathop{\max}\limits_{{b_{l,m}}}$ $\ \mathop{\min}\limits_{k\in{\cal K}}\ {{R_{k}}}$ (49)
${\rm{s.t.}}$ $\ {b_{l,m}}\in{\tilde{\cal C}}$,
$\ {\left|{{b_{l,m}}-{b_{l,m^{\prime}}}}\right|^{2}}\geq\tilde{d}_{\min}^{2},\ \forall m^{\prime}\neq m.$